# Yuang Li, Yinglu Li, Min Zhang, Chang Su, Mengxin Ren, Xiaosong Qiao, Xiaofeng Zhao, Mengyao Piao, Jiawei Yu, Xinglin Lv, Miaomiao Ma, Yanqing Zhao, Hao Yang

### v3

@@ Abstract
End-to-end automatic speech recognition (ASR) systems often struggle to recognize rare name entities, such as personal names, organizations, and terminologies not frequently encountered in the training data. This paper presents Contextual Biasing Whisper (CB-Whisper), a novel ASR system based on OpenAI's Whisper model that can recognize user-defined name entities by performing open-vocabulary keyword-spotting (OV-KWS) using the hidden states of Whisper encoder. The recognized entities are used as prompts for the Whisper decoder. We first propose a multitask training approach with OV-KWS and ASR tasks to optimize the model. Experiments show that this approach substantially improves the entity recalls compared to the original Whisper model on Chinese Aishell hot word subsets and two internal code-switch test sets. However, we observed a slight increase in mixed-error-rate (MER) on internal test sets due to catastrophic forgetting. To address this problem and use different sizes of the Whisper model without finetuning, we propose to use OV-KWS as a separate module and construct a spoken form prompt to prevent hallucination. The OV-KWS module consistently improves MER and Entity Recall for whisper-small, medium, and large models.
Index Terms: speech recognition, contextual biasing, keyword-spotting
@@ 1 Introduction
End-to-end (E2E) automatic speech recognition (ASR) models [1, 2, 3] have gained popularity in recent years for their simplicity and unified architecture. However, they suffer from low recall of proper nouns that are rare in the training data. Contextual biasing (CB) is a technique that leverages external knowledge of the expected words in the speech input to mitigate the long-tail word problem. One of the early approaches is shallow fusion [4, 5], which builds an n-gram finite state transducer from a predefined list of hot words and boosts their scores during beam search decoding. However, this approach has some drawbacks, such as the challenge of finding the optimal fusion weight. Therefore, researchers have proposed deep fusion methods [6, 7, 8, 9, 10], which train a contextual encoder together with the ASR model from scratch. These methods embed the entity words and the speech signal into the same feature space, and the decoder uses both contextual and acoustic information to generate the transcription. To adapt existing ASR models, some methods use adaptors [11, 12] to modify the intermediate features of the ASR model or pointer networks [13] to modify the output distributions. Moreover, shallow and deep fusion methods can be combined to further enhance the performance [14, 15].
@@
OpenAI's Whisper [16] is a state-of-the-art ASR model based on the Transformer [17] architecture and it was trained on a large-scale dataset of 680k hours of speech data. Unlike previous models that require a separate biasing module, it can adaptively bias the generation of entity words by providing a simple prompt to the decoder. In this paper, we propose a novel method called contextual biasing Whisper (CB-Whisper) which incorporates an open-vocabulary keyword-spotting (OV-KWS) module between the encoder and the decoder of the Whisper to search entity words before constructing prompts. The OV-KWS module is inspired by vision-based KWS methods [18, 19] and uses features from the Whisper encoder to construct a cosine similarity matrix. To train the OV-KWS model and enhance Whisper's ability to understand prompts, we introduce a multitask training approach with OV-KWS and ASR tasks. Our system enhances entity recalls with absolute improvements of up to 80% on Aishell hot word subsets [20] and up to 10% on Internal code-switch datasets. However, we noticed that CB-Whisper finetuned on a small-scale dataset has higher mixed-error-rates (MERs) compared to the Whisper model on real-world Internal datasets due to catastrophic forgetting. Therefore, we propose to freeze the Whisper model and use OV-KWS as an independent module. To align the format of the prompt with the original Whisper training setup, we devised a spoken form prompt that mimics the history transcription. For whisper-small, medium, and large models, our system can achieve absolute MER reductions of 2.4%, 1.4%, and 1.8% respectively on Internal datasets. Note that the OV-KWS is a lightweight module using the Whisper-small encoder, making it computationally efficient.
@@
Figure 1: (a) Entity Features are extracted using TTS followed by Whisper encoder. (b) The flowchart of the CB-Whisper model. (c) Illustration of OV-KWS module.
@@ 2.1 System overview
Before the system can be deployed, a database is created by extracting acoustic representations for pre-defined entity words. As shown in Figure 1 (a), the speech signal for each entity word is generated with a Text-to-Speech (TTS) model. Then, the whisper encoder is used to obtain hidden states that will be later matched with the features of the input utterance.
@@
As illustrated in Figure 1 (b), the proposed CB-Whisper identifies entity words before the decoder by utilizing the hidden states of the encoder. In detail, we first obtain the hidden representations of the input utterance by summing up the hidden states from all Whisper encoder layers with a learnable weight for each layer. Next, we calculate the cosine similarity matrix between the input utterance features and the stored features of each entity word. A binary Convolutional Neural Network (CNN) classifier then determines if the entity word is present in the utterance by recognizing the diagonal pattern in the similarity matrix (Figure 1 (c)). Finally, the predicted entity words are used as prompts to guide the whisper decoder so that these words can be recognized more accurately.
2.2 Open-vocabulary keyword-spotting
In the proposed system, OV-KWS is formulated as a binary classification task. A deep neural network predicts the presence of an entity word in the utterance by taking the cosine similarity matrix between the hidden states of entity words and the input utterance as input. As shown in Figure 1 (c), if an entity word exists in the utterance, a diagonal pattern can be observed. Since such a pattern is a local feature, we choose CNN as the classification model. The usage of the similarity matrix rather than the original hidden states improves the model's generalisability and allows the detection of words that were not seen during training. Following the common approach to using multi-level self-supervised features for downstream tasks [21], we use the weighted sum of multi-layer features from Whisper-encoder instead of a single-layer feature.
The OV-KWS module has a linear computational complexity to the number of keywords. To enhance the efficiency, we employed a very lightweight CNN network with only four layers and 0.2 million parameters. Furthermore, multiple keywords can be batched and executed in parallel on GPU.
An easy way to construct the dataset for the CNN classifier is to select positive words according to the transcription and to choose negative words randomly. However, this approach is too simplistic, as the negative samples may have significant differences from the positive samples, resulting in overfitting. Thus, we adopted hard negative samples [22], which are more likely to have overlaps with positive samples, increasing the difficulty of the classification task.
2.3 Multitask finetuning
@@
The CNN classifier's capacity is limited by its small size. Therefore, it is necessary to finetune the Whisper model for the OV-KWS task while preserving its ASR ability. Hence, we propose to finetune all parameters of the Whisper model on the OV-KWS task and the ASR task jointly with the loss function shown in Equation 1. Additionally, during finetuning, we incorporate words from the ground-truth transcription into the prompt following the naive prompt format in Table 1. This enables the Whisper decoder to better understand the prompt format and leverage the word in the prompt.
@@ 2.4 Prompting Whisper decoder
Prompt is a term that refers to a segment of text that provides the context, or the objective for the generation of large language models. Whisper, an encoder-decoder model, also supports prompting because the decoder can be viewed as an internal language model. The decoder used the transcription of the previous utterance as a prompt during training to incorporate contextual information for the recognition of the following utterance. During inference, any text-only prompt that is relevant to the input utterance can serve as historical context for decoding.
Table 1: Prompts for whisper decoder.
naive prompt
实体1, 实体2, 实体3
entity 1, entity 2, entity 3
spoken form prompt
今天演讲的主题是这个呃,实体1、实体2、实体3。好,那我就继续讲。
The topic of today’s speech is, ah, entity 1, entity 2, entity 3. Okay, then I’ll continue.
For the multitask training experiments, we adopt the naive prompt format, which consists of the plain concatenation of all the entity words. This format works well for the in-domain test sets (Aishell), but it may induce hallucination on more difficult spontaneous talks that involve code-switching situations. Moreover, the finetuned model’s ASR performance may slightly degrade, as it was finetuned on a small-scale dataset. To address this issue and integrate the OV-KWS module into different versions of the Whisper model without training, we propose an alternative to multitask finetuning, which is to use OV-KWS as an independent module and keep the Whisper model fixed. To prevent hallucination, we designed the spoken form prompt (Table 1) which incorporates disfluencies and indicated that the subsequent speech was a spontaneous talk. Compared to the naive prompt, the spoken form prompts have a more similar form to the history transcriptions used during training of the Whisper model.
@@ 3.1 Training configurations
For the multitask training experiments, we selected the Whisper-small model. In subsequent experiments, we also compared the Whisper-small, Whisper-medium, and Whisper-large-v2 models without finetuning, using OV-KWS as a plug-in.
For the binary CNN classifier, we employed a very lightweight architecture with four layers. Each layer had 128, 128, 256, and 256 channels respectively, and a kernel size of three. We adopted a two-phase multitask training scheduler. In the first phase, we froze the Whisper encoder and trained only the CNN classifier for OV-KWS and the Whisper decoder for ASR, for 10 epochs, with a batch size of 128 and a learning rate of $5×10^{-5}$. The total number of trainable parameters was 153.8 million (CNN 0.2 million + Decoder 153.6 million). In the second phase, we unfroze the Whisper encoder and trained the entire model for another 50 epochs, with a learning rate of $3×10^{-5}$. The number of trainable parameters increased to 241.9 million. Note that for the ASR task, we applied a random concatenation strategy [23] to increase the length of the utterance. We performed all the training on four NVIDIA T4 GPUs. During inference, we used a beam size of five for beamsearch decoding.
@@ 3.2 Datasets
To construct the training set for OV-KWS and ASR, we utilized three datasets: Aishell-1 [24], LibriSpeech [25], and TALCS [26]. Aishell-1 is a Mandarin ASR dataset with 150 hours of speech recordings. LibriSpeech is an English ASR dataset, from which we selected the train-clean subset containing 100 hours of speech recordings. TALCS is a Mandarin-English codeswitch dataset with 578 hours of speech data from English teaching scenes. For each utterance in these datasets, we randomly sampled positive words and generated speech using edge-tts  [^1] https://github.com/rany2/edge-tts. We compared the performance of the model trained with only Chinese data (zh) and the model trained with the mixture of three datasets (zh + en).
@@
3.3 Evaluation matrics
We evaluated the performance of OV-KWS using the F1 score, which is the harmonic mean of the precision and recall. For ASR, we adopted two metrics: mixed-error-rate (MER) and entity recall. MER is a modified version of character-error-rate (CER) that can handle code-switching situations between English and Chinese. In this metric, we treat each Chinese character and each English word as a single unit. Therefore, MER is equivalent to CER for Chinese and word-error-rate (WER) for English.
@@ 4.1 Results for open-vocabulary keyword spotting
We evaluated the performance of the CNN classifier for OV-KWS using the F1 score metric. Table 3 shows the results of different training settings. We first trained the model with a frozen Whisper encoder on the Aishell dataset only and obtained F1 scores of around 0.6 on the Aishell test sets and 0.68 on the internal datasets. Then, we evaluated the model trained with a mixture of three datasets and observed consistent improvement in the F1 scores. Finally, we unfroze the Whisper encoder, which increased the number of parameters of the OV-KWS task from 0.2 to 88.4 million and achieved significant improvement in the F1 scores. The model trained with the Aishell dataset performed slightly better on the Aishell dev and test sets, with F1 scores of 0.85 and 0.89, respectively. On the internal datasets, the model trained with the mixed dataset achieved higher F1 scores, reaching around 0.9. The comparison between single-task training and multitask training is presented in Table 4. The results indicate that jointly optimizing ASR and OV-KWS does not adversely affect the performance of the OV-KWS task and even leads to a higher F1 score on the Internal-1 dataset. Figure 2 illustrates the weight assigned to each layer of the Whisper encoder. The hidden states from the 8th to the 12th layers have a higher impact on the OV-KWS task. Moreover, the weight for the deeper layers increased slightly after finetuning, suggesting that the Whisper encoder was more adapted to the OV-KWS task.
Table 3: The performance of OV-KWS module measured by F1 score. We compared the models trained with Chinese-only data (zh) and Chinese-English-codeswitch data (zh + en). We also compared the results with the Whisper model frozen or unfrozen.
@@
Table 4: The comparison between single task training and multitask training of OV-KWS model measured by F1 score.
@@
Table 5: The performance of CB-Whisper measured by MER (%) / Entity Recall (%). The Whisper-small model was unfrozen and trained jointly with the OV-KWS module. "zh" means the model was finetuned on Chinese-only data. "zh+en" represents the model was trained on Chinese-English-codeswitch data. We prompted the Whisper decoder with ground-truth entities and predicted entities respectively.
Model Whisper-small (baseline) CB-Whisper-small (zh) CB-Whisper-small (zh + en)
@@ 4.2 Results for CB-Whisper
The multitask training of our CB-Whisper model resulted in the ASR performance shown in Table 5. Compared to the original Whisper-small model, the finetuned CB-Whisper-small model achieved significantly lower MER and higher Entity recall on the Aishell-dev and test sets. The model that was trained on the mixed dataset exhibited higher entity recall across all datasets, indicating better prompting function and generalisability. Remarkably, the CB-Whisper model with predicted entities surpassed the Whisper-small model with ground-truth entities in entity recall, reaching 88.4% and 87.7% on the Aishell-dev and test sets respectively. Furthermore, the CB-Whisper model outperformed the state-of-the-art SeACo-Paraformer model [20], which has entity recalls of 86% and 87% on the Aishell dev and test sets.
@@
However, we found that though entity recall could be enhanced on internal test sets, the MERs increased slightly. This was mainly because we fine-tuned CB-Whisper on small-scale datasets, and the domain of the training data (audiobook and teaching scenes) differed from that of the test sets (technical talks). As a result, Whisper’s generalisability was reduced. Therefore, we propose a different approach, which involves freezing the original Whisper model and using OV-KWS as a separate module. As shown in Table 6, we used the Whisper-small model in two ways: either with a CNN classifier and the original whisper-small's features (the first block) or with an OV-KWS model (the second block). The second option requires more computation, as two different whisper-encoders were used, but it can achieve higher entity recalls. A benefit of this method is that the OV-KWS model can be applied to various Whisper models without fine-tuning. However, we found that with the naive prompt, the MER can increase compared to the original model. The main reason was that the Whisper model tended to omit filler words and disfluencies if the prompt was in a well-formatted text form. This problem could be alleviated by using spoken form prompts that mimicked a presentation transcription. Compared to the naive prompt and no prompt, spoken form prompts consistently improved the MER. As an example, using the OV-KWS model and the spoken form prompt, the MER decreased from 4.4% to 2.9% and from 5.6% to 3.6% on the Internal-1 and Internal-2 subsets respectively for the Whisper-large-v2 model  [^2] We have conducted experiments with the Whisper-large-v3 model, but it did not outperform the v2 model.. In addition, the proposed method attained entity recalls of 97.7% and 96.9% for the two datasets, which are considerably superior to the original model’s performance of 87.3% and 79.9%, respectively.
@@
Table 6: The performance of various Whisper models with OV-KWS as an independent module measured by MER (%) / Entity Recall (%). We compared the naive prompt and the spoken form prompt.
Model Prompt Internal-1 Internal-2
CB-Whisper-small × 5.9 / 78.5 9.7 / 60.8
(frozen) Naive 5.5 / 91.5 8.5 / 87.9
Spoken 4.2 / 91.2 6.6 / 89.0
Whisper-small Naive 5.4 / 94.0 9.0 / 90.2
+ OV-KWS Spoken 4.0 / 95.9 6.8 / 91.1
Whisper-medium × 4.8 / 84.2 6.7 / 70.4
+ OV-KWS Naive 6.8 / 94.3 14.6 / 90.3
Spoken 4.0 / 95.6 4.8 / 94.4
Whisper-large-v2 × 4.4 / 87.3 5.6 / 79.9
+ OV-KWS Naive 4.7 / 97.4 6.6 / 95.1
Spoken 2.9 / 97.7 3.6 / 96.9

### v4

# Miaomiao Ma, Daimeng Wei, Shimin Tao, Hao Yang

@@ Abstract
The recognition of rare named entities, such as personal names and terminologies, is challenging for automatic speech recognition (ASR) systems, especially when they are not frequently observed in the training data. In this paper, we introduce keyword spotting enhanced Whisper (KWS-Whisper), a novel ASR system that leverages the Whisper model and performs open-vocabulary keyword spotting (OV-KWS) on the hidden states of the Whisper encoder to recognize user-defined named entities. These entities serve as prompts for the Whisper decoder. To optimize the model, we propose a multitask training approach that learns OV-KWS and contextual-ASR tasks. We evaluate our approach on Chinese Aishell hot word subsets and two internal code-switching test sets and show that it significantly improves the entity recall compared to the original Whisper model. Moreover, we demonstrate that the OV-KWS can be a plug-and-play module to enhance the ASR error correction methods and frozen Whisper models.
keywords:
speech recognition, keyword spotting, prompt
@@ 1 Introduction
Recent years have witnessed the rise of end-to-end (E2E) automatic speech recognition (ASR) models [1, 2, 3] due to their simplicity and unified architecture. However, these models struggle to recognize proper nouns that infrequently occur in the training data. A common approach is to construct a list of hot words and apply it to shallow fusion [4, 5] or deep fusion [6, 7, 8, 9, 10]. Shallow fusion enhances the hot words’ score during beam search decoding, but it requires tuning the optimal fusion weight. The deep fusion method trains a contextual encoder jointly with the ASR model from scratch. These methods map the entity words and the speech signal into the same feature space, and the decoder leverages both contextual and acoustic information to generate the transcription. To adapt existing ASR models, some methods use adaptors [11, 12] to modify the intermediate features of the ASR model or pointer networks [13] to revise the output distributions. Moreover, the shallow and deep fusion methods can be combined to enhance the performance [14, 15].
@@
Whisper [16] stands as a cutting-edge ASR model. It leverages the Transformer architecture [17] and was trained on an extensive dataset comprising 680k hours of speech data. Unlike its predecessors, which necessitate a separate biasing module, Whisper dynamically adjusts its output by incorporating a straightforward prompt during decoding. In this paper, we introduce an innovative technique called KWS-Whisper (Keyword Spotting Enhanced Whisper). This method integrates an open-vocabulary keyword spotting (OV-KWS) module between the encoder and decoder of Whisper. The OV-KWS module draws inspiration from vision-based keyword spotting methods [18, 19] and constructs a cosine similarity matrix using features from the Whisper encoder. To enhance Whisper’s comprehension of prompts, we employ a multitask training approach that combines OV-KWS and contextual-ASR tasks. Our system enhances entity recall with absolute improvements of up to 80% on Aishell hot word subsets [20] and up to 10% on internal code-switching datasets. However, we noticed that the KWS-Whisper fine-tuned on a small-scale dataset has higher mixed error rates (MERs) than the Whisper model on real-world internal datasets due to catastrophic forgetting. Therefore, we propose to use OV-KWS as an independent module for frozen Whisper models. We conducted experiments for various scenarios: 1) ASR error correction based on Pinyin or the large language model (LLM); 2) directly prompting the Whisper decoder with spoken-form prompts that resemble the historical transcription. The results show that our system can achieve significant MER reductions of 2.4%, 1.4%, and 1.8% on internal datasets for Whisper-small, medium, and large models respectively. Compared to our previous work [21], the lightweight OV-KWS module and the multitask training approach markedly enhance the system’s adaptability, efficiency, and accuracy.
@@
Figure 1: (a) Entity Features are extracted using TTS followed by Whisper encoder. (b) Flowchart of the KWS-Whisper model. (c) Illustration of the OV-KWS module.
@@ 2.1 System overview
A database is first constructed by extracting acoustic representations for pre-defined entity words. As shown in Figure 1 (a), a text-to-speech (TTS) model followed by the Whisper encoder is used to extract hidden states, which are later matched with the input utterance features. As illustrated in Figure 1 (b), the proposed KWS-Whisper identifies entity words before decoding by leveraging encoder hidden states. Specifically, we aggregate hidden representations from all Whisper encoder layers, weighted by learnable coefficients. A cosine similarity matrix is then computed between the input utterance features and stored entity word features. A binary Convolutional Neural Network (CNN) detects the presence of an entity word by recognizing the diagonal pattern in the similarity matrix (Figure 1 (c)). Finally, predicted entity words guide the Whisper decoder for improved recognition accuracy.
2.2 Open-vocabulary keyword spotting
OV-KWS is framed as a binary classification task. The model predicts the presence of an entity word within an utterance by analyzing the cosine similarity matrix between the hidden states of entity words and the input utterance. Notably, if an entity word exists in the utterance, a diagonal pattern emerges (Figure 1 (c)). Given that this is a local pattern, we opt for a CNN as our classification model. To extract features, we follow a common approach of using the weighted sum of multi-layer features from the Whisper encoder [22]. The OV-KWS module exhibits linear computational complexity to the number of keywords. For higher efficiency, we employ an ultra-lightweight CNN comprising only four layers and 0.2 million parameters. Additionally, we leverage GPU parallelization to batch and execute multiple keywords simultaneously. An easy way to construct the dataset for the CNN classifier is to select positive words according to the transcription and choose negative words randomly. However, this approach is too simplistic, as the negative samples may have significant differences from the positive samples, resulting in overfitting. As such, we adopt hard negative samples [23], which are more likely to have overlaps with positive samples, increasing the difficulty of the classification task.
2.3 Multitask fine-tuning
@@
The CNN classifier’s capacity is limited by its small size. Therefore, it is necessary to fine-tune the Whisper model for the OV-KWS task while preserving its ASR ability. To this end, we propose to fine-tune all parameters of the Whisper model on both the OV-KWS task and the contextual-ASR task jointly with the loss function shown in Equation 1. During fine-tuning, we incorporate words from the ground-truth transcription into the prompt following the naive prompt format in Table 1. This enables the Whisper decoder to better understand the prompt format and leverage the word in the prompt.
@@
{CJK}
UTF8gbsn
Table 1: Prompts for the Whisper decoder.
naive prompt 实体1, 实体2, 实体3
(translation) entity 1, entity 2, entity 3
spoken-form prompt 今天演讲的主题是这个呃,实体1、实体2、实体3。好,那我就继续讲。
(translation) The topic of today’s speech is, ah, entity 1, entity 2, entity 3. Okay, then I’ll continue.
@@ 2.4 Prompting Whisper decoder
During training, the Whisper decoder incorporates contextual information by using the transcription of the previous utterance as a prompt. In inference, any relevant text-only prompt can serve as the historical context for decoding. For multitask training, we adopt a straightforward prompt format (i.e., concatenating all entity words) which performs remarkably well on in-domain test sets (Aishell). However, this format can lead to hallucinations in more complex spontaneous conversations involving code-switching. Additionally, the fine-tuned model’s ASR performance may slightly degrade due to its small-scale training dataset. Therefore, we explore using OV-KWS as a plug-in for correcting ASR errors and prompting frozen Whisper models.
2.5 OV-KWS plug-in for frozen Whisper models
We can use the OV-KWS plug-in for:
  • ASR error correction: 1) Pinyin-based method: We segment each transcription into words and replace the Chinese word if it has the same Pinyin (pronunciation) as the entity word. For English words, we compute the edit distance and replace the word if the distance is below a pre-defined threshold. 2) LLM-based method: We use ChatGLM3-6B [24] and the prompt in Table 2. In our experiments, we provide two examples for the LLM.
  • Prompting Whisper decoder: Since the Whisper models are not fine-tuned, we use spoken-form prompt (Table 1) to prevent hallucinations which incorporates disfluencies and indicate that the subsequent speech was a spontaneous talk. Compared to the naive prompt, the spoken-form prompt has a more similar form to the historical transcriptions used during the training of the Whisper model.
{CJK}
UTF8gbsn
Table 2: The prompt for LLM-based ASR error correction. The translation of the Chinese prompt is given in parentheses with italics.
User:
请根据可能出现的关键词修改ASR文本。
(Please modify the ASR transcription according to possible keywords.)
关键词 (keywords):{word1},{word2}。
识别文本 (recognized text):{ASR transcription}
Response:
修改文本 (modified text):
@@ 3.1 Training configurations
We selected the Whisper-small model for the multitask training. In subsequent experiments, we also compared the Whisper-small, medium, and large-v2 versions without fine-tuning. For the CNN, we employed a lightweight four-layer architecture. Each layer had 128, 128, 256, and 256 channels respectively, and a kernel size of three. In the first training phase, we froze the Whisper encoder and trained the CNN classifier for OV-KWS and the Whisper decoder for ASR, for 10 epochs, with a batch size of 128 and a learning rate of $5×10^{-5}$. The total number of trainable parameters was 153.8 million (CNN 0.2 million + decoder 153.6 million). In the second training phase, we unfroze the Whisper encoder and trained the entire model for another 50 epochs, with a learning rate of $3×10^{-5}$. The number of trainable parameters increased to 241.9 million. Note that for the ASR task, we applied a random concatenation strategy [25] to increase the length of the utterance. We used a beam size of five for decoding.
@@ 3.2 Datasets
To construct the training set for OV-KWS and ASR, we utilized three datasets: Aishell-1 [26], LibriSpeech [27], and TALCS [28]. Aishell-1 is a Mandarin ASR dataset with 150 hours of speech recordings. LibriSpeech is an English ASR dataset, from which we selected the train-clean subset containing 100 hours of speech recordings. TALCS is a Mandarin-English code-switching dataset with 578 hours of speech data from English teaching scenes. For each utterance in these datasets, we randomly sampled positive words and generated speech using edge-tts  [^1] https://github.com/rany2/edge-tts. We compared the performance of the model trained with only Chinese data (zh) and the model trained with the mixture of three datasets (zh + en).
@@
3.3 Evaluation metrics
We evaluated the performance of OV-KWS using the F1 score, which is the harmonic mean of the precision and recall. For ASR, we adopted two metrics: mixed error rate (MER) and entity recall. MER is a modified version of character error rate (CER) that can handle code-switching situations between English and Chinese. In this metric, we treat each Chinese character and each English word as a single unit. Therefore, MER is equivalent to CER for Chinese and word error rate (WER) for English.
@@ 4.1 Results for open-vocabulary keyword spotting
We evaluated the performance of the CNN classifier for OV-KWS using the F1 score. Table 4 shows the results of different training settings. We first trained the model with a frozen Whisper encoder on the Aishell dataset only and obtained F1 scores of around 0.6 on the Aishell test sets and 0.68 on the internal datasets. Then, we evaluated the model trained with a mixture of three datasets and observed consistent improvement in the F1 scores. Finally, we unfroze the Whisper encoder, which increased the number of parameters of the OV-KWS task from 0.2 million to 88.4 million and achieved significant improvement in the F1 scores. The model trained with the Aishell dataset performed slightly better on the Aishell-dev and Aishell-test sets, with F1 scores of 0.85 and 0.89, respectively. On the internal datasets, the model trained with the mixed dataset achieved higher F1 scores, reaching around 0.9. The comparison between single-task training and multitask training is presented in Table 5. The results indicate that simultaneous optimization of ASR and OV-KWS does not adversely affect the performance of the OV-KWS task instead, it leads to a higher F1 score on the Internal-1 dataset. Figure 2 illustrates the weight assigned to each layer of the Whisper encoder. The hidden states from the 8th to the 12th layers have a greater impact on the OV-KWS task. Moreover, the weight for the deeper layers increased slightly after fine-tuning, suggesting that the Whisper encoder is more adapted to the OV-KWS task.
Table 4: The performance of the OV-KWS module measured by F1 score. We compared the models trained with Chinese-only data (zh) and Chinese-English-code-switching data (zh + en). We also compared the results with the Whisper model frozen or unfrozen.
@@
Table 5: The comparison between single-task training and multitask training of the OV-KWS model measured by F1 score.
@@
Table 6: The performance of KWS-Whisper measured by MER (%) / Entity Recall (%). The Whisper-small model was unfrozen and trained jointly with the OV-KWS module. ”zh” means the model was fine-tuned on Chinese-only data. ”zh + en” represents the model was trained on Chinese-English-code-switching data. We prompted the Whisper decoder with ground-truth entities and predicted entities respectively.
Model Whisper-small (baseline) KWS-Whisper-small (zh) KWS-Whisper-small (zh + en)
@@ 4.2 Results for KWS-Whisper
The multitask training of our KWS-Whisper model yielded the ASR performance shown in Table 6. Compared to the original Whisper-small model, the fine-tuned KWS-Whisper model achieved a significantly lower MER and higher entity recall on the Aishell-dev and Aishell-test sets. The model that was trained on the mixed dataset exhibited higher entity recall across all datasets, indicating better prompting function and generalizability. Remarkably, the KWS-Whisper model with predicted entities surpassed the Whisper-small model with ground-truth entities in entity recall, reaching 88.4% and 87.7% on the Aishell-dev and test sets respectively. Furthermore, the KWS-Whisper model outperformed the state-of-the-art SeACo-Paraformer model [20], which has entity recalls of 86% and 87% on the Aishell-dev and Aishell-test sets.
@@
However, we found that though the entity recalls could be enhanced on internal test sets, the MERs increased slightly. This was mainly because we fine-tuned KWS-Whisper on small-scale datasets, and the domain of the training data (audiobook and teaching scenes) differed from that of the test sets (technical talks). As a result, Whisper’s generalizability was reduced. Hence, we suggest using OV-KWS as a plug-in for ASR error correction or applying frozen Whisper models. Table 7 shows that error correction with Chinese Pinyin, using the recognized entity from the OV-KWS model, consistently improved the MER and entity recalls for all Whisper model sizes. The LLM-based correction achieved higher entity recalls than the naive Pinyin-based method, as it incorporated contextual information. However, the MER increased due to the LLM’s hallucinations. We also employed the prompting approach in KWS-Whisper, without fine-tuning various Whisper models. This method outperformed error correction–based methods significantly, as it considered acoustic features. Nevertheless, we discovered that the naive prompt slightly raised the MER compared to the original model. The main reason was that the Whisper model tends to omit filler words and disfluencies if the prompt is in well-formatted text. This problem could be alleviated by using the spoken-form prompt that mimics a presentation transcription. Compared to the naive prompt and no prompt, spoken-form prompts consistently improved the MER. For example, using the OV-KWS model and the spoken-form prompt, the MER decreased from 4.4% to 2.9% and from 5.6% to 3.6% on the Internal-1 and Internal-2 subsets respectively for the Whisper-large-v2 model. In addition, the best system attained entity recalls of 97.7% and 96.9% for the two datasets, considerably superior to the original model’s performance of 87.3% and 79.9%, respectively.
@@
Table 7: The performance of various systems with OV-KWS as an independent module measured by MER (%) / Entity Recall (%). We investigated the effects of two ASR error correction techniques: Pinyin-based and LLM-based, and two types of prompts for the Whisper decoder: naive and spoken-form.
Model Method Internal-1 Internal-2
Whisper- × 5.9 / 78.5 9.7 / 60.8
small Correction-Pinyin 5.7 / 82.9 9.1 / 68.2
Correction-LLM 6.9 / 85.2 12.7 / 74.3
Prompt-Naive 5.4 / 94.0 9.0 / 90.2
Prompt-Spoken 4.0 / 95.9 6.8 / 91.1
Whisper × 4.8 / 84.2 6.7 / 70.4
-medium Correction-Pinyin 4.8 / 85.2 6.3 / 78.3
Correction-LLM 6.1 / 88.9 9.3 / 80.9
Prompt-Naive 6.8 / 94.3 14.6 / 90.3
Prompt-Spoken 4.0 / 95.6 4.8 / 94.4
Whisper × 4.4 / 87.3 5.6 / 79.9
-large-v2 Correction-Pinyin 4.3 / 88.9 5.5 / 83.9
Correction-LLM 6.0 / 89.9 7.7 / 85.1
Prompt-Naive 4.7 / 97.4 6.6 / 95.1
Prompt-Spoken 2.9 / 97.7 3.6 / 96.9